# Around the Merino–Welsh conjecture: improving Jackson’s inequality

### v1

@@ 4. The main lemma
In this section we give the main ingredient of the proofs of Theorems 1.3 and 1.4.
@@
Let $H=(A,B,E)$ be a bipartite graph. For $x\geq 2$ and a fixed $0\leq s\leq 1$ let us introduce the function
@@
$\widetilde{T}_{H}(x,0)\geq\prod_{v\in A}\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\cdot\prod_{u\in B}\left(s-\frac{s}{d_{u}+1}\prod_{v\in N(u)}\gamma_{x,s}(d_{v})\right).$
@@
The first idea is that we only take those values for which $x_{u}=1-y_{u}\leq s$ for $u\in B$. Hence
$\widetilde{T}_{H}(x,0)\geq\int_{[0,1]^{A}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
For each $v\in A$ we decompose the integral over $[0,1]$ to $[0,s]$ and $[s,1]$, this way we decomposed the integral $[0,1]^{A}$ to the sum of $2^{|A|}$ integrals. Since $x_{u}\leq s$ for $u\in B$ we immediately get that if $x_{v}\geq s$, then $X_{v}(\underline{x},\underline{y})=x$. Let $T\subseteq A$ be the set of those vertices $v\in A$ for which $x_{v}\geq s$. Then
@@
$=\sum_{T\subseteq A}(x(1-s))^{|T|}\int_{[0,s]^{A\setminus T}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
Let $N(T)\subseteq B$ be the neighbors of $T$. If $u\in N(T)$, then $\widehat{Y}_{u}(\underline{x},\underline{y}):=Y_{u}(\underline{x},\underline{y})=1$, because the vertex $u$ cannot be active anymore. Hence
$I_{T}:=\int_{[0,s]^{A\setminus T}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}$
corresponds to $s^{|A\setminus T|+|B|}$ times a permutation Tutte polynomial, where an active vertex $i\in A\setminus T$ gets a weight $x^{(i)}=x$ while in $B$ an active vertex $j$ gets a weight $y^{(j)}=1$ if it is $N(T)$, and $y^{(j)}=0$ if it is $B\setminus N(T)$, so for this term we have a lower bound from Lemma 2.13:
@@
$\geq s^{|A\setminus T|+|B|}\prod_{i\in A}\mathbb{E}[\widehat{X}_{i}]\cdot\prod_{j\in B}\mathbb{E}[\widehat{Y}_{j}]$
@@
$\widetilde{T}_{H}(x,0)$ $=\sum_{T\subseteq A}(x(1-s))^{|T|}I_{T}$
$\geq\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot s^{|A\setminus T|+|B|}\prod_{v\in A\setminus T}\left(1+\frac{x-1}{d_{v}+1}\right)\prod_{u\in B\setminus N(T)}\left(1-\frac{1}{d_{u}+1}\right).$
@@
Note that $\mu$ is a product measure $\mu=\bigoplus_{v\in A}\mu_{v}$, where
@@
For each $u\in B$ consider the function
@@
$\widetilde{T}_{H}(x,0)$ $\geq\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot\prod_{v\in A\setminus T}\left(\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s\right)\cdot\prod_{u\in B\setminus N(T)}\frac{d_{u}s}{d_{u}+1}\cdot\prod_{u\in N(T)}s$
@@
All functions $f_{u}$ are monotone increasing on the space $\{0,1\}^{A}$ identified with the subsets $T$ of $A$, since adding vertices to $T$ can only increase $N(T)$, so by Harris’ inequality we have
$Z\sum_{T\subseteq A}\mu(T)\prod_{u\in B}f_{u}(T)\geq Z\prod_{u\in B}\left(\sum_{T\subseteq A}\mu(T)f_{u}(T)\right).$
@@
$\widetilde{T}_{H}(x,0)\geq\prod_{v\in A}\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\cdot\prod_{u\in B}\left(s-\frac{s}{d_{u}+1}\prod_{v\in N(u)}\gamma_{x,s}(d_{v})\right).$
@@
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\prod_{v\in V}\left[\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(d_{u})\right)\right].$
@@
If $x\geq 1$ and $0\leq s\leq 1$, then the function $\gamma_{x,s}(d)$ is monotone decreasing in $d$. Furthermore, $\gamma_{x,s}(d)\leq s$.
@@
$sx-sx(x(1-s)+s)=sx(1-x)(1-s)\leq 0$
which is clearly true. The second statement follows from $d+x\leq(d+1)x$ as follows:
$\gamma_{x,s}(d)=\frac{(d+x)s}{(d+x)s+(d+1)x(1-s)}\leq\frac{(d+x)s}{(d+x)s+(d+x)(1-s)}=s.$
Definition 4.4.
Let
For any $10\leq\gamma<$ 1 we have
$G(\infty,x,s):=\lim_{d\to\infty}G(d,x,s,\gamma)=(s+x(1-s))s.$
@@
Let $H$ be a bipartite graph with minimum degree at least $\delta$. Then for $x\geq 2$ and $0\leq s\leq 1$ we have
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(\delta))\geq\prod_{v\in V}G(d_{v},x,s,s).$
@@
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)$ $\geq\prod_{v\in V}\left[\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(d_{u})\right)\right]$
$\geq\prod_{v\in V}\left[\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(\delta)\right)\right]$
@@
Let $x\geq 2$ and $\gamma\leq s<1$. A sufficient condition for
$G(d,x,s,\gamma)\geq G(d+1,x,s,\gamma)$
is $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}\geq\gamma.$
Furthermore, the function $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}$ is monotone increasing in $d$ for $x\geq 1$ and $d\geq 2$, so it is enough to check the above inequality for some fixed $d_{0}$ to conclude that the sequence $G(d,x,s,\gamma)$ is decreasing for $d\geq d_{0}$.
@@
We have
$G(d,x,s,\gamma)-G(d+1,x,s,\gamma)$
$=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma^{d}\right)-\left(\frac{(d+1+x)s}{d+2}+x(1-s)\right)\left(s-\frac{s}{d+2}\gamma^{d+1}\right)$
$=s^{2}\left(\frac{d+x}{d+1}-\frac{d+1+x}{d+2}\right)-\frac{s}{d+1}\gamma^{d}\left(\frac{(d+x)s}{d+1}+x(1-s)\right)+\frac{s}{d+2}\gamma^{d+1}\left(\frac{(d+1+x)s}{d+2}+x(1-s)\right)$
$\geq\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{s}{d+1}\gamma^{d}\left(\frac{(d+x)s}{d+1}+x(1-s)\right)$
$\geq\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{sx}{d+1}\gamma^{d}$
which is non-negative if $\frac{x-1}{x(d+2)}\geq\gamma^{d-1}$. Finally, the derivative of $h(t):=\frac{1}{t-1}\ln\left(\frac{x-1}{x(t+2)}\right)$ is
@@
$h^{\prime}(t)\geq\frac{\ln(t+2)}{(t-1)^{2}}-\frac{1}{(t-1)(t+2)}\geq\frac{\ln(3)}{(t-1)^{2}}-\frac{1}{(t-1)(t+2)}>0$
@@
In this section we prove that
$T_{M}(x,0)T_{M}(0,x)\geq T_{M}(1,1)^{2}$
whenever $x\geq 2.355$. By the transfer lemma it is enough to prove that
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\widetilde{T}_{H}(1,1)^{2}$
for any bipartite graph $H$ with minimum degree at least $1$. Note that $\widetilde{T}_{H}(1,1)=1$.
@@
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(1)).$
Since $\widetilde{T}_{H}(1,1)^{2}=1$ it is enough if $G(d,x,s,\gamma_{x,s}(1))\geq 1$ for all $d\geq 1$. Let
@@
These numbers are chosen in such a way that $G(1,x,s,\gamma_{x,s}(1))=1$. It is easy to check that for $d=9$ we have $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}\geq s\geq\gamma_{x,s}(1)$. By Lemma 4.6 this implies that for $d\geq 9$ we have $G(d,x,s,\gamma_{x,s}(1))\geq G(\infty,x,s)>1$. It is also easy to check (see the jupyter notebook or Table 1 in the Appendix) that $G(d,x,s,\gamma_{x,s}(1))\geq 1$ for $d\in\{2,\dots,10\}$. So with $x=\frac{3+\sqrt{5}}{2}$ we have
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\widetilde{T}_{H}(1,1)^{2}$
@@
Idea 2. Observe that we can assume that $H$ is connected since $\widetilde{T}_{H}(x,y)=\prod_{i=1}^{k}\widetilde{T}_{H_{i}}(x,y)$ if $H$ has connected components $H_{1},\dots,H_{k}$. The statement $\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq 1$ is trivially true for $H=K_{2}$, the complete graph on $2$ vertices. So we can assume that $H$ has at least $3$ vertices. On the other hand, if $H$ is connected and has at least $3$ vertices, then a degree one vertex cannot be adjacent to another degree one vertex. Thus it is enough if
$G(1,x,s,\gamma_{x,s}(2))\geq 1\ \ \text{and}\ \ G(d,x,s,\gamma_{x,s}(1))\geq 1$
for all $d\geq 2$. This is satisfied if $x=2.54$ and $s=0.76$. (See Table 2.)
Idea 3. Observe that it is not really necessary that $G(1,x,s,\gamma)\geq 1$, it is quite enough if for a non-leaf vertex the product with its leaf neighbors is at least $1$. If $x\geq 2.2$, then any vertex can have at most $2$ leaf neighbors, since for the star $S_{4}$ on $4$ vertices we have
$\frac{1}{x}\widetilde{T}_{S_{4}}(x,0)\widetilde{T}_{S_{4}}(0,x)=\frac{1}{x}\cdot\frac{x^{3}+x^{2}+x}{4}\cdot\frac{x}{4}=\frac{x^{3}+x^{2}+x}{16}>1$
@@
so by the gluing lemma (Lemma 2.14) we know that a graph containing a pending $S_{4}$ cannot be a minimal counter example. So it is enough to check if for every $d\geq 2$ we have
$G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{\min(2,d-1)}\geq 1$

### v2

@@ 4. The main lemma
In this section, we give the main ingredient of the proofs of Theorems 1.3 and 1.4.
@@
Let $H=(A,B,E)$ be a bipartite graph. For $x\geqslant 2$ and a fixed $0\leqslant s\leqslant 1$ let us introduce the function
@@
$\widetilde{T}_{H}(x,0)\geqslant\prod_{v\in A}\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\cdot\prod_{u\in B}\left(s-\frac{s}{d_{u}+1}\prod_{v\in N(u)}\gamma_{x,s}(d_{v})\right).$
@@
The first idea is that we only consider those $((x_{v})_{v\in A},(y_{u})_{u\in B})$ for which $x_{u}=1-y_{u}\leqslant s$ for $u\in B$. Hence
$\widetilde{T}_{H}(x,0)\geqslant\int_{[0,1]^{A}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
For each $v\in A$ we decompose the integral over $[0,1]$ to $[0,s]$ and $[s,1]$, this way we decomposed the integral $0,1]^{A}$ to the sum of $2^{|A|}$ integrals. Since $x_{u}\leqslant s$ for $u\in B$ we immediately get that if $x_{v}\geqslant s$, then $v$ is active and so $X_{v}(\underline{x},\underline{y})=x$. Let $T\subseteq A$ be the set of those vertices $v\in A$ for which $x_{v}\geqslant s$. Then
@@
$=\sum_{T\subseteq A}(x(1-s))^{|T|}\int_{[0,s]^{A\setminus T}}\int_{[1-s,1]^{B}}\prod_{v\in A\setminus T}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
Let $N(T)\subseteq B$ be the neighbors of $T$. If $u\in N(T)$, then the vertex $u$ cannot be active anymore, and so let us define $\widehat{Y}_{u}(\underline{x},\underline{y}):=Y_{u}(\underline{x},\underline{y})=1$. Hence
$I_{T}:=\int_{[0,s]^{A\setminus T}}\int_{[1-s,1]^{B}}\prod_{v\in A\setminus T}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}$
corresponds to $s^{|A\setminus T|+|B|}$ times a permutation Tutte polynomial on $(A\setminus T,B)$, where an active vertex $i\in A\setminus T$ gets a weight $x^{(i)}=x$ while in $B$ an active vertex $j$ gets a weight $y^{(j)}=1$ if it is $N(T)$, and $y^{(j)}=0$ if it is $B\setminus N(T)$, so for this term we have a lower bound from Lemma 2.12:
@@
$\geqslant s^{|A\setminus T|+|B|}\prod_{i\in A}\mathbb{E}[\widehat{X}_{i}]\cdot\prod_{j\in B}\mathbb{E}[\widehat{Y}_{j}]$
@@
$\widetilde{T}_{H}(x,0)$ $\geqslant\sum_{T\subseteq A}(x(1-s))^{|T|}I_{T}$
$\geqslant\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot s^{|A\setminus T|+|B|}\prod_{v\in A\setminus T}\left(1+\frac{x-1}{d_{v}+1}\right)\prod_{u\in B\setminus N(T)}\left(1-\frac{1}{d_{u}+1}\right).$
@@
Note that $\mu$ is a product measure $\mu=\bigotimes_{v\in A}\mu_{v}$, where
@@
For each $u\in B$ consider the function $f_{u}:\{0,1\}^{A}\to\mathbb{R}$ defined as
@@
$\widetilde{T}_{H}(x,0)$ $\geqslant\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot\prod_{v\in A\setminus T}\left(\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s\right)\cdot\prod_{u\in B\setminus N(T)}\frac{d_{u}s}{d_{u}+1}\cdot\prod_{u\in N(T)}s$
@@
All functions $f_{u}$ are monotone increasing on the space $\{0,1\}^{A}$ identified with the subsets $T$ of $A$, since adding vertices to $T$ can only increase $N(T)$, so by Harris’s inequality (Lemma 2.9) we have
$Z\sum_{T\subseteq A}\mu(T)\prod_{u\in B}f_{u}(T)\geqslant Z\prod_{u\in B}\left(\sum_{T\subseteq A}\mu(T)f_{u}(T)\right).$
@@
$\widetilde{T}_{H}(x,0)\geqslant\prod_{v\in A}\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\cdot\prod_{u\in B}\left(s-\frac{s}{d_{u}+1}\prod_{v\in N(u)}\gamma_{x,s}(d_{v})\right).$
@@
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\prod_{v\in V}\left[\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(d_{u})\right)\right].$
Definition 4.3.
Let us introduce the notation
$\mathbb{G}\left(d,x,s,\{\gamma_{j}\}_{1\leqslant j\leqslant d}\right):=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\prod_{j=1}^{d}\gamma_{j}\right).$
Furthermore, let
For any $0\leqslant\gamma<1$ we have
$G(\infty,x,s):=\lim_{d\to\infty}G(d,x,s,\gamma)=(s+x(1-s))s.$
Clearly, with this notation we can rewrite Corollary 4.2 as
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right).$
Next, we prove some monotonicity properties of the function $\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$.
@@
If $x\geqslant 1$ and $0\leqslant s\leqslant 1$, then the function $\gamma_{x,s}(d)$ is monotone decreasing in $d$. Furthermore, $\gamma_{x,s}(d)\leqslant s$.
@@
$sx-sx(x(1-s)+s)=sx(1-x)(1-s)\leqslant 0$
which is clearly true. The second statement follows from $d+x\leqslant(d+1)x$ as follows:
$\gamma_{x,s}(d)=\frac{(d+x)s}{(d+x)s+(d+1)x(1-s)}\leqslant\frac{(d+x)s}{(d+x)s+(d+x)(1-s)}=s.$
@@
Let $H$ be a bipartite graph with minimum degree at least $\delta$. Then for $x\geqslant 2$ and $0\leqslant s\leqslant 1$ we have
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(\delta))\geqslant\prod_{v\in V}G(d_{v},x,s,s).$
@@
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)$ $\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
$\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(\delta)\}_{u\in N(v)}\right)$
@@
Let $x\geqslant 1$ and $0\leqslant\gamma\leqslant s<1$. If $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}\geqslant\gamma$, then
$G(d,x,s,\gamma)\geqslant G(d+1,x,s,0),$
consequently
$G(d,x,s,\gamma)\geqslant G(d+1,x,s,0)\geqslant G(d+1,x,s,\gamma).$
Furthermore, the function $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}$ is monotone increasing in $d$ for $x\geqslant 1$ and $d\geqslant 2$, so it is enough to check the above inequality for some fixed $d_{0}$ to conclude that the sequence $G(d,x,s,\gamma)$ is decreasing for $d\geqslant d_{0}$.
@@
The inequality $G(d+1,x,s,0)\geqslant G(d+1,x,s,\gamma)$ is trivial. We only need to prove the inequality $G(d,x,s,\gamma)\geqslant G(d+1,x,s,0)$. We have
$G(d,x,s,\gamma)-G(d+1,x,s,0)$
$=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma^{d}\right)-\left(\frac{(d+1+x)s}{d+2}+x(1-s)\right)s$
$=s^{2}\left(\frac{d+x}{d+1}-\frac{d+1+x}{d+2}\right)-\frac{s}{d+1}\gamma^{d}\left(\frac{(d+x)s}{d+1}+x(1-s)\right)$
$=\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{s}{d+1}\gamma^{d}\left(\frac{(d+x)s}{d+1}+x(1-s)\right)$
$\geqslant\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{sx}{d+1}\gamma^{d}$
$\geqslant\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{sx}{d+1}s\gamma^{d-1}$
$=\frac{xs^{2}}{d+1}\left(\frac{x-1}{x(d+2)}-\gamma^{d-1}\right)$
which is non-negative if $\frac{x-1}{x(d+2)}\geqslant\gamma^{d-1}$. Finally, the derivative of $h(t):=\frac{1}{t-1}\ln\left(\frac{x-1}{x(t+2)}\right)$ is
@@
$h^{\prime}(t)\geqslant\frac{\ln(t+2)}{(t-1)^{2}}-\frac{1}{(t-1)(t+2)}\geqslant\frac{\ln(3)}{(t-1)^{2}}-\frac{1}{(t-1)(t+2)}>0$
@@
In this section, we prove that
$T_{M}(x,0)T_{M}(0,x)\geqslant T_{M}(1,1)^{2}$
whenever $x\geqslant 2.355$. By the transfer lemma it is enough to prove that
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\widetilde{T}_{H}(1,1)^{2}$
for any bipartite graph $H$ with minimum degree at least $1$. Recall that $\widetilde{T}_{H}(1,1)=1$.
@@
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(1)).$
Since $\widetilde{T}_{H}(1,1)^{2}=1$ it is enough if $G(d,x,s,\gamma_{x,s}(1))\geqslant 1$ for all $d\geqslant 1$. Let
@@
These numbers are chosen in such a way that $G(1,x,s,\gamma_{x,s}(1))=1$. It is easy to check that for $d_{0}=9$ we have $\left(\frac{x-1}{x(d_{0}+2)}\right)^{1/(d_{0}-1)}\geqslant s\geqslant\gamma_{x,s}(1)$. By Lemma 4.4 this implies that for $d\geqslant d_{0}$ we have $G(d,x,s,\gamma_{x,s}(1))\geqslant G(\infty,x,s)>1$. It is also easy to check (see the jupyter notebook or Table 1 in the Appendix) that $G(d,x,s,\gamma_{x,s}(1))\geqslant 1$ for $d\in\{2,\dots,8\}$. So with $x=\frac{3+\sqrt{5}}{2}$ we have
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\widetilde{T}_{H}(1,1)^{2}$
@@
Idea 2. Observe that we can assume that $H$ is connected since $\widetilde{T}_{H}(x,y)=\prod_{i=1}^{k}\widetilde{T}_{H_{i}}(x,y)$ if $H$ has connected components $H_{1},\dots,H_{k}$. The statement $\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant 1$ is trivially true for $H=K_{2}$, the complete graph on $2$ vertices. So we can assume that $H$ has at least $3$ vertices. On the other hand, if $H$ is connected and has at least $3$ vertices, then a degree one vertex cannot be adjacent to another degree one vertex. We claim that it is enough to check that if
(5.1) $G(1,x,s,\gamma_{x,s}(2))\geqslant 1\ \ \text{and}\ \ G(d,x,s,\gamma_{x,s}(1))\geqslant 1$
for all $d\geqslant 2$. Indeed,
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)$ $\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
$=\prod_{v\in V:d_{v}=1}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)\cdot\prod_{v\in V:d_{v}\geqslant 2}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
$\geqslant\prod_{v\in V:d_{v}=1}G(1,x,s,\gamma_{x,s}(2))\cdot\prod_{v\in V:d_{v}\geqslant 2}G(d_{v},x,s,\gamma_{x,s}(1))$
$\geqslant 1.$
We claim that the conditions of 5.1 are satisfied if $x=2.54$ and $s=0.76$, see Table 2. Clearly, the condition $G(1,x,s,\gamma_{x,s}(2))\geqslant 1$ is just a matter of computation. To prove that $G(d,x,s,\gamma_{x,s}(1))\geqslant 1$ for all $d\geqslant d_{0}=9$, we use the same argument as before: for $d\geqslant d_{0}$ the sequence is $G(d,x,s,\gamma_{x,s}(1))$ is monotone decreasing by Lemma 4.6 and checking that
$\left(\frac{x-1}{x(d_{0}+2)}\right)^{1/(d_{0}-1)}>\gamma_{x,s}(1).$
This gives that
$G(d,x,s,\gamma_{x,s}(1))\geqslant G(\infty,x,s)>1$
for all $d\geqslant d_{0}$. Since $G(d,x,s,\gamma_{x,s}(1))>1$ for $2\leqslant d\leqslant 8$ we get this inequality for all $d\geqslant 2$.
Idea 3. The next idea is to group the degree $1$ vertices together with their neighbors:
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)$ $\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
$=\prod_{v\in V:d_{v}\geqslant 2}\left(\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)\cdot\prod_{u\in N(v):d_{u}=1}\mathbb{G}\left(d_{u},x,s,\{\gamma_{x,s}(d_{u^{\prime}})\}_{u^{\prime}\in N(u)}\right)\right)$
$=\prod_{v\in V:d_{v}\geqslant 2}\left(\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)\cdot\prod_{u\in N(v):d_{u}=1}G\left(1,x,s,\gamma_{x,s}(d_{v})\right)\right)$
$\geqslant\prod_{v\in V:d_{v}\geqslant 2}\left(\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(1)\}_{u\in N(v)}\right)\cdot\prod_{u\in N(v):d_{u}=1}G\left(1,x,s,\gamma_{x,s}(d_{v})\right)\right)$
$=\prod_{v\in V:d_{v}\geqslant 2}\left(G(d_{v},x,s,\gamma_{x,s}(1))\prod_{u\in N(v):d_{u}=1}G\left(1,x,s,\gamma_{x,s}(d_{v})\right)\right)$
It would be enough for us if for all $v$ with $d_{v}\geqslant 2$ we have
$G(d_{v},x,s,\gamma_{x,s}(1))\prod_{u\in N(v):d_{u}=1}G\left(1,x,s,\gamma_{x,s}(d_{v})\right)\geqslant 1.$
In order to prove such an inequality, we need an upper bound for the number of degree $1$ neighbors of a vertex $v$. We claim that if $x\geqslant 2.2$, then any vertex can have at most $2$ leaf neighbors. To prove this, we use the gluing lemma. First, observe that for the star $S_{4}$ on $4$ vertices we have
$\frac{1}{x}\widetilde{T}_{S_{4}}(x,0)\widetilde{T}_{S_{4}}(0,x)=\frac{1}{x}\cdot\frac{x^{3}+x^{2}+x}{4}\cdot\frac{x}{4}=\frac{x^{3}+x^{2}+x}{16}>1,$
@@
so by the gluing lemma (Lemma 2.13) we know that a graph containing a pendant $S_{4}$ cannot be a minimal counter example. This means that in a minimal counter example every vertex has at most $2$ neighbors of degree $1$. So, it is enough to check if for every $d\geqslant 2$ we have
$\min_{0\leqslant k\leqslant\min(2,d-1)}G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{k}\geqslant 1.$
Here we used that if $d=2$ then it cannot have two neighbors of degree $1$, since then $K_{1,2}$ would be a connected component of $H$, but it is clearly not a counter example. We show that for $x=2.36$ and $s=0.78$ these inequalities are satisfied. We can check this inequality separately for $d=2$. For $d>2$ we have
$\ \min_{0\leqslant k\leqslant\min(2,d-1)}G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{k}$
$=\min(G(d,x,s,\gamma_{x,s}(1)),G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{2}).$
(Clearly, if $G(1,x,s,\gamma_{x,s}(d))>1$ then the minimum is achieved at the first term, otherwise the minimum is achieved at the second term.)
In order to check that the above minimum is at least $1$ we proceed as before. Suppose that $G(\infty,x,s)\geqslant 1,$ and for some $d_{0}$ we have
(5.2) $G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{2}\geqslant 1,$
and for $d\geqslant d_{0}$ the sequence $G(d,x,s,\gamma_{x,s}(1))$ is monotone decreasing. To check this latter property, we only need to check
$\left(\frac{x-1}{x(d_{0}+2)}\right)^{1/(d_{0}-1)}>\gamma_{x,s}(1).$
Then for $d\geqslant d_{0}$ we have
$G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{2}\geqslant G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{2}\geqslant 1,$